for what values of a and b lim x?8 ((x²+1)÷(x+1))-ax-b =0

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
\lim_{x\rightarrow \infty }\frac{x^{2}+1}{x+1}-ax-b = 0
\lim_{x\rightarrow \infty }\frac{x^{2}+1-ax^{2}-bx-ax-b}{x+1}= 0
\lim_{x\rightarrow \infty }\frac{x^{2}(1-a)+x(-a-b)+1-b}{x+1}= 0
1-a = 0
a = 1
-a-b = 0
b = -a = -1
a = 1, b = -1
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty